Each content area in Mathematics contributes towards the acquisition of specific skills. One 

focus area emphasised is that “the learner should recognize and describe properties of 

numbers and operations, including (i) identity properties, (ii) factors, (iii) multiples, and (iv) 

commutative; (vi) associative and (vii) distributive properties” Describe in brief, what the 

identity property entails and provide two relevant examples. (Hint: use different basic 

operations to clarify your answer)


1
Expert's answer
2021-06-16T14:30:54-0400

Identity property entails a set numbers with unique properties. It applies to a group of number in a set as it does not apply to a number only.

Example

Identity property of addition

When we add zero to a number, the sum is always the number, the sum is always the number.

"(n+0=n)\\2+0=2"

Identity property of multiplication when a number is multiplied by 2. The product is the number itself "(n\times 1=n)\\2\times1=2."